# M2D2: Exploring General-purpose Audio-Language Representations Beyond CLAP

### v1

@@ V Conclusion
This study pursued a general-purpose audio-language representation that provides both general-purpose audio and CLAP features to serve diverse audio applications. For this purpose, we proposed M2D2, a new method combining an SSL M2D and CLAP in a two-stage pre-training and utilizing a recent LLM-based sentence encoder. In the first stage, M2D2 learns a generalizable audio feature using the rich semantic features of the LLM-based sentence encoder, and in the second stage, it learns a CLAP feature using the learned audio feature. While M2D learns highly generalized audio features, the powerful LLM-based embeddings help learn a better CLAP feature through the audio feature. Experiments validated the effectiveness of M2D2 as a general-purpose audio-language representation, highlighted by the SOTA performance of audio features for AudioSet fine-tuning and music tasks and top performance of CLAP features in zero-shot classification, audio-text retrieval, and audio captioning. We conducted ablation studies to analyze M2D2 pre-training and found that two-stage training strategy design and managing the effect of the CLAP objective are vital for learning effective features. We have made our models available online and hope they help the progress of diverse audio studies101010To reviewers, we will update our repository after the notification..

### v2

@@ V Conclusion
This study pursued a general-purpose audio-language representation that provides both general-purpose audio and CLAP features to serve diverse audio applications. For this purpose, we proposed M2D-CLAP, a novel method combining an SSL M2D and CLAP in a multi-stage pre-training and utilizing a recent LLM-based sentence encoder. In the first stage, M2D-CLAP learns a generalizable audio feature using the rich semantic features of the LLM-based sentence embedding model, and in the second stage, it learns a CLAP feature using the learned audio feature. It further refines features by fine-tuning. Experiments validated the effectiveness of M2D-CLAP as a general-purpose audio-language representation. Its audio features performed well, particularly in AudioSet fine-tuning and music tasks, and its CLAP features performed well in zero-shot classification, audio-text retrieval, and audio captioning. These results demonstrate the novelty of M2D-CLAP in providing both high-performing general audio and CLAP features. We conducted ablation studies to analyze M2D-CLAP and found that a multi-stage training strategy design and managing the effect of the CLAP objective are vital for learning effective features. We release our models to advance diverse audio studies.